# Open Materials Generation with Stochastic Interpolants

### v1

@@ VI Discussion and Conclusions
During hyperparameter optimization for CSP models, we observed a tradeoff between match rate and RMSE—which is only computed if matched—between the ideal and generated crystals. By adjusting the relative cost between the position, cell, and species terms in the loss, we observe that the main learning challenge lies in the accurate prediction of the atomic coordinates, which always have a significantly higher relative weight in calculating the full loss function in Eq. 19 in the appendix. Unsurprisingly, we find that as the number of generated structures matching known compositions in the test dataset increases, the structural fidelity of these matches—quantified by the accompanying RMSE—also tends to increase. We quantify this tradeoff in Fig. 7 in the appendix. However, match rate remains the primary metric of interest, as the generated structure can always be relaxed after generation if seeking minimum energy structures (i.e., at zero temperature).
For the CSP task, OMG significantly outperforms previous approaches on all datasets. We highlight the particularly strong performance of the trigonometric interpolant in achieving a high match rate for the perov-5 dataset as shown in Tab. 1. Unlike other datasets, perov-5 has a fixed number of atoms $N$ per unit cell and a fixed (cubic) cell with varying side lengths and identical fractional positions. In contrast, in other datasets, no unique representation of the periodic repeat unit is imposed on flows, meaning the model does not learn the invariance (or even equivariance) to the choice of periodic repeat unit. [^8] Using Niggli reduction during learning to enforce a unique choice of unit cell on structures from our datasets is not sufficient for enforcing this invariance during generation of structures. This likely contributes to the difficulty of unconstrained flow-based models in generating highly symmetric structures. The perov-5 dataset presents a unique case where this invariance does not need to be learned, making it an ideal benchmark for evaluating which positional interpolant can best learn structural relationship when the species are fixed. The match rate achieved using the trigonometric interpolant in Tab. 1 surpasses (by 1.6×1.6$\times$1.6 ×) our best-performing linear interpolant model for perov-5, which has a lower match rate of 51.86% (without filtering invalid atoms) and an RMSE of 0.0757 (state-of-the-art for this metric). It is possible that the superior performance of the trigonometric interpolant arises from its ability to generate more circuitous flow trajectories compared to the strictly geodesic paths imposed by the linear interpolant, akin to the reasoning behind using latent variables to enhance learning in SIs [33].
Finally, we note trends in Tab. 4 among the best-performing models identified through hyperparameter optimization. In particular, we observe a tradeoff between the choice of positional interpolant and the optimal level of ‘species noise’ $\eta$ which sets the probability that an atom will change its identity if already in an unmasked state (see Appendix A.5), with linear interpolants favoring higher species noise. Additionally, we find a correlation between enabling element order permutation (to minimimize the distance between $x_{0}$ and $x_{1}$) and the use of geodesic (linear) interpolation paths, with linear interpolants favoring minimum distance permutation. Therefore, we conclude that DNG models built on linear interpolants likely benefit from targeted tuning strategies that are unnecessary for non-geodesic positional interpolants.
Table 4: Optimally performing DNG trials from OMG. All models are trained with an ODE sampling scheme and ordered by decreasing species noise $\eta$. The values of whether the element ordering is permuted to minimize traveled distance is also included.
Positional Min. dist. Species
Interpolant permutation noise
Linear, $\gamma=0$ True 35.46
Linear, $\gamma=0$ True 31.29
Linear, $\gamma=0$ True 7.08
SBD, $\gamma=0$ False 1.95
EncDec, $\gamma=\sin^{2}{\pi t}$ False 0.85
Trig, $\gamma=\sqrt{1.19t(1-t)}$ True 0.81
In conclusion, we demonstrate OMG’s capabilities for learning the manifold of stable materials and generating novel structures. We show state-of-the-art performance of SIs across nearly all benchmarks for both crystal structure prediction as well as de novo generation of stable, novel, and unique structures. In our discussion, we highlight several ways in which hyperparameters and interpolation paths influence learning, though a more systematic theoretical analysis is needed to fully understand why specific choices affect model performance. Future work can also extend the flexibility of OMG to additional interpolating functions. We underscore the importance of flexible ML frameworks like OMG, which can adapt to different types of materials datasets by optimizing the generative model accordingly. Our work represents a key step forward in applications of machine-learning methods to materials discovery.
@@ VII Acknowledgements
This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. The authors gratefully acknowledge the support of compute resources provided by the Empire AI Consortium. The authors also acknowledge UFIT Research Computing (http://www.rc.ufl.edu) at the University of Florida, the Minnesota Supercomputing Institution (MSI) (http://www.msi.umn.edu) at the University of Minnesota, and resources provided by Drexel University through NSF Grant OAC-2320600, for providing computational resources and support that have contributed to the research results reported in this publication. The authors acknowledge funding from NSF Grant OAC-2311632. P.H. also acknowledges support by a grant from the Simons Foundation (839534, MT).
Appendix A Implementation details of stochastic interpolants
A.1 OMG Framework
@@
Figures 4 and 5 summarize the training and the integration pipeline of the OMG framework, respectively. Depending on the specific task, there are several stochastic interpolants at once. For CSP, one stochastic interpolant considers lattice vectors $\bm{L}$, and another one considers fractional coordinates $\bm{X}$. The model output of CSPNet (see Appendix B.1) depends on the full structural representation $\{\bm{A},\bm{X},\bm{L}\}$ and time $t$, where $\bm{A}$ are the atomic species. For the DNG task, we additionally use discrete flow matching for the atomic species $\bm{A}$ [42].
@@
During the numerical integration in the CSP task, $\bm{X}$ and $\bm{L}$ are integrated jointly while $\bm{A}$ is fixed. For DNG, $\bm{A}$ is evolved according to discrete flow matching [42] (see Appendix A.5). For the SDE sampling scheme in Fig. 5, one chooses a time-dependent noise $\varepsilon(t)$ that only appears during integration and not during training. Also, $\gamma(t)$ has to be unequal zero in order to prevent the divergence in $1/\gamma(t)$. However, since $\gamma(t)$ necessarily vanishes at times $t=0$ and $t=1$ (see Appendix A.2), one should choose a time-varying $\varepsilon(t)$ that vanishes near these endpoints [33].
@@
Figure 4: Training pipeline of the OMG framework: A batch of structures is drawn from a dataset with target distribution $\rho_{1}$. Every structure $x_{1}\sim\rho_{1}$ is connected with a structure $x_{0}$ from the base distribution $\rho_{0}$ with stochastic interpolants that yield the interpolated structure $x_{t}=x(t,x_{0},x_{1},z)$ and the drift $b_{t}=\partial_{t}x_{t}$ at time $t\sim\mathcal{U}(0,1)$, possibly using a random variable $z\sim\mathcal{N}(0,\bm{I})$. The model CSPNet predicts $b^{\theta}_{t}=b^{\theta}(t,x_{t})$ and $z^{\theta}_{t}=z(t,x_{t})$ and its parameters are minimized based on the MSE losses in Eqs 2 and 3.
@@
Figure 5: Numerical integration pipeline of the OMG framework: An initial structure $x_{0}$ from the base distribution $\rho_{0}$ is numerically integrated following either an ODE or an SDE based on the model predictions $b_{t}^{\theta}$ and $z_{t}^{\theta}$. For an SDE, one can choose a noise $\varepsilon(t)$ during integration.
A.2 Interpolant choice
@@
We are concerned with interpolants with the form specified in Eq. 1. Here, the following conditions must be met [33]:
$\alpha(0)=\beta(1)=1,\quad\alpha(1)=\beta(0)=\gamma(0)=\gamma(1)=0,\quad\gamma(t)>0\,\forall t\in(0,1)$ (5)
Under these constraints, the form of the SI is relatively flexible, and many different interpolants can be defined (see Tab. 5). An additional advantage of SI is that the base distribution can be arbitrary as in CFM [39]. We find that particular forms of the stochastic interpolants result in better performance than others. To add to the flexibility of the SI framework, ODE or SDE integration can be used to sample from a trained SI model. We thoroughly explore the use of both of these techniques to investigate the performance of SI on different elements of the crystal unit cell.
@@
Arbitrary $\rho_{0}$ trig $\cos\frac{\pi}{2}t$ $\sin\frac{\pi}{2}t$ $\sqrt{at(1-t)}$
@@
linear $1-t$ $t$ 00
Gaussian $\rho_{0}$ trig $\cos\frac{\pi}{2}t$ $\sin\frac{\pi}{2}t$ 00
SBD $\sqrt{1-t^{2}}$ $t$ 00
Mirror 00 $1$ $\sqrt{at(1-t)}$
A.3 Antithetic sampling
As shown by Albergo et al. [33], inspection of possible $\gamma$ terms and of Eq. 2 reveals that the loss function can become unstable around $t=0$ and $t=1$. To account for this, we implement antithetic sampling. This requires to compute the loss at both $x^{+}$ and $x^{-}$ where:
@@ A.3 Antithetic sampling
A.4 Interpolation with periodic boundary conditions
We show in Fig. 6 how we implement periodic versions of interpolants in order to represent fractional coordinates in a unit cell with periodic boundary conditions. We emphasize that this procedure is important not only for the choice of interpolant, but also for the addition of the latent variable $\gamma(t)z$ which also moves the interpolation trajectory away from the geodesic.
@@ A.4 Interpolation with periodic boundary conditions
A.5 DFM Details
DFM allows for generative modeling of discrete sequences of tokens while respecting the discrete nature of the design space. As discussed, a parameterized neural network $p_{1|t}^{\theta}(x_{1}|x_{t})$ is learned, which attempts to predict the final sequence from the current sequence. Borrowing from Campbell et al. [42], we choose a conditional rate matrix $R_{t}(x_{t},i|x_{1})$ generating the conditional flow $p_{t|1}(x_{t}|x_{1})$ of the form:
@@ A.5 DFM Details
In practice, Campbell et al. [42] show that we need not compute a full expectation, but rather, simply draw $x_{1}\sim p_{1|t}^{\theta}(x_{1}|x_{t})$, evaluate the conditional rate matrix, and perform an update of $x_{t}$ to $x_{t+\Delta t}$ with discrete time step $\Delta t$ directly from this by sampling $x_{t+\Delta t}$ according to
@@
A.6 Permuting particle order to achieve minimum interpolated distance
We experiment with biasing the initial distribution such that we learn interpolant paths that are shorter on average than independently chosen $(x_{0},x_{1})$ pairs. Formally, our coupling is conditional on the sampled $(x_{0},x_{1})$ and is defined as

### v2

@@ VI Discussion and Conclusions
We adapt stochastic interpolants (SIs) for material generation tasks and propose Open Materials Generation (OMatG), a material-generation framework that unifies score-based diffusion and conditional flow‑matching approaches under the umbrella of SIs. By incorporating an equivariant graph representation of crystal structures and explicitly handling periodic boundary conditions, OMatG jointly models spatial coordinates, lattice vectors, and discrete atomic species in a cohesive flow-based pipeline. Our extensive experiments on crystal structure prediction and de novo generation tasks demonstrate that OMatG sets a new state of the art in generative modeling for inorganic materials discovery, yielding more stable, novel, and unique structures than either pure diffusion or pure conditional flow‑matching counterparts. We underscore the importance of flexible ML frameworks like OMatG, which can adapt to different types of materials datasets by optimizing the generative model accordingly. Our work represents a key step forward in applications of machine-learning methods to materials discovery. Looking forward, we plan to extend the flexibility of OMatG to additional interpolating functions, improve on the evaluation metrics and datasets, and investigate how different SIs influence the discovery of suitable materials.
@@ VII Acknowledgements
The authors thank Shenglong Wang at NYU IT High Performance Computing and Gregory Wolfe for their resourcefulness and valuable support. The authors acknowledge funding from NSF Grant OAC-2311632. P. H. and S. M. also acknowledge support from the Simons Center for Computational Physical Chemistry (Simons Foundation grant 839534, MT). The authors gratefully acknowledge use of the research computing resources of the Empire AI Consortium, Inc, with support from the State of New York, the Simons Foundation, and the Secunda Family Foundation. Moreover, the authors gratefully acknowledge the additional computational resources and consultation support that have contributed to the research results reported in this publication, provided by: IT High Performance Computing at New York University; the Minnesota Supercomputing Institute (http://www.msi.umn.edu) at the University of Minnesota; UFIT Research Computing (http://www.rc.ufl.edu) and the NVIDIA AI Technology Center at the University of Florida in part through the AI and Complex Computational Research Award; Drexel University through NSF Grant OAC-2320600.
Appendix A Acronyms
Table 4 provides a list of acronyms used throughout this paper for reference.
Table 4: Acronym definitions.
Acronym Full Name
AIRSS Ab initio Random Structure Searching
CDVAE Crystal Diffusion Variational Autoencoder
CFP Chemical Formula Prediction
CFM Conditional Flow Matching
CTMC Continuous‐Time Markov Chain
CSP Crystal Structure Prediction
DFT Density Functional Theory
DFM Discrete Flow Matching
DNG De Novo Generation
EGNN Equivariant Graph Neural Network
GNoME Graph Networks for Materials Exploration
MLIP Machine-Learned Interatomic Potential
NequIP Neural Equivariant Interatomic Potentials
ODE Ordinary Differential Equation
OMatG Open Materials Generation
RFM Riemannian Flow Matching
SBD Score‐Based Diffusion
SBDM Score‐Based Diffusion Model
SDE Stochastic Differential Equation
SI Stochastic Interpolant
SMACT Semiconducting Materials from Analogy and Chemical Theory
SUN Stable, Unique, and Novel
Appendix B Implementation Details of Stochastic Interpolants
B.1 OMatG Framework
@@
Figures 4 and 5 summarize the training and the integration pipeline of the OMatG framework, respectively. Depending on the specific task, there are several stochastic interpolants at once. For CSP, one stochastic interpolant considers lattice vectors $\bm{L}$, and another one considers fractional coordinates $\bm{X}$. The model output of CSPNet (see Appendix C.1) depends on the full structural representation $\{\bm{A},\bm{X},\bm{L}\}$ and time $t$, where $\bm{A}$ are the atomic species. For the DNG task, we additionally use discrete flow matching for the atomic species $\bm{A}$ [44].
@@
During the numerical integration in the CSP task, $\bm{X}$ and $\bm{L}$ are integrated jointly while $\bm{A}$ is fixed. For DNG, $\bm{A}$ is evolved according to discrete flow matching [44] (see Appendix B.6). For the SDE sampling scheme in Fig. 5, one chooses a time-dependent noise $\varepsilon(t)$ that only appears during integration and not during training (see Appendix B.4). Also, $\gamma(t)$ has to be unequal zero in order to prevent the divergence in $1/\gamma(t)$. However, since $\gamma(t)$ necessarily vanishes at times $t=0$ and $t=1$ (see Appendix B.2), one should choose a time-varying $\varepsilon(t)$ that vanishes near these endpoints (see Appendix B.4[33].
@@
Figure 4: Training pipeline of the OMatG framework: A batch of structures is drawn from a dataset with target distribution $\rho_{1}$. Every structure $x_{1}\sim\rho_{1}$ is connected with a structure $x_{0}$ from the base distribution $\rho_{0}$ with stochastic interpolants that yield the interpolated structure $x_{t}=x(t,x_{0},x_{1},z)$ and the drift $b_{t}=\partial_{t}x_{t}$ at time $t\sim\mathcal{U}(0,1)$, possibly using a random variable $z\sim\mathcal{N}(0,\bm{I})$. The model CSPNet predicts $b^{\theta}_{t}=b^{\theta}(t,x_{t})$ and $z^{\theta}_{t}=z(t,x_{t})$ and its parameters are minimized based on the MSE losses in Eqs (2) and (3) [see also Eq. (32)].
@@
Figure 5: Numerical integration pipeline of the OMatG framework: An initial structure $x_{0}$ from the base distribution $\rho_{0}$ is numerically integrated following either an ODE or an SDE based on the model predictions $b_{t}^{\theta}$ and $z_{t}^{\theta}$. For an SDE, one can choose a noise $\varepsilon(t)$ during integration.
B.2 Interpolant Choice
@@
In this work, we are concerned with spatially linear interpolants of the form specified in Eq. (1). The following conditions must be met [33]:
$\alpha(0)=\beta(1)=1,\quad\alpha(1)=\beta(0)=\gamma(0)=\gamma(1)=0,\quad\gamma(t)>0~{}\forall t\in(0,1).$ (5)
Under these constraints, the form of the SI is relatively flexible, and many different interpolants can be defined. Also, the base distribution can be arbitrary as in CFM [38, 40, 41]. In this work, we mostly rely on interpolants originally defined in Albergo et al. [33] and listed in Tab. 5. For the VP SBD interpolant, we allow for different widths $\sigma_{0}$ of the Gaussian base distribution $\rho_{0}$.
In Appendix B.9, we introduce additional VP and VE SBD interpolants beyond the one in Tab. 5 that are implemented in OMatG. The encoder-decoder (enc-dec) interpolant as defined in Tab. 5 evolves samples from the base distribution $\rho_{0}$ to follow an intermediate Gaussian distribution with variance $1$ at the switch time $T_{\text{switch}}=0.5$, before mapping them to a sample from $\rho_{1}$. This can be generalized to arbitrary variances $a>0$ and switch times $T_{\text{switch}}\in(0,1)$:
$\begin{split}\alpha(t)&=\cos^{2}\left(\frac{\pi(t-T_{\text{switch}}t)^{p}}{(T_{\text{switch}}-T_{\text{switch}}t)^{p}+(t-T_{\text{switch}}t)^{p}}\right)\,1_{[0,T_{\text{switch}})}(t),\\ \beta(t)&=\cos^{2}\left(\frac{\pi(t-T_{\text{switch}}t)^{p}}{(T_{\text{switch}}-T_{\text{switch}}t)^{p}+(t-T_{\text{switch}}t)^{p}}\right)\,1_{(T_{\text{switch}},1]}(t),\\ \gamma(t)&=\sqrt{a}\sin^{2}\left(\frac{\pi(t-T_{\text{switch}}t)^{p}}{(T_{\text{switch}}-T_{\text{switch}}t)^{p}+(t-T_{\text{switch}}t)^{p}}\right),\end{split}$ (6)
where $p\geq 1/2$. We consider the cases $p\in\{1/2,1\}$ and note that the general interpolant in Eq. (6) reduces to the interpolant in Tab. 5 for $a=1$, $p=1$, and $T_{\text{switch}}=0.5$.
@@
Arbitrary $\rho_{0}$ trig $\cos\left(\frac{\pi}{2}t\right)$ $\sin\left(\frac{\pi}{2}t\right)$ $\sqrt{at(1-t)}$
@@
Gaussian $\rho_{0}$ VP SBD $\sqrt{1-t^{2}}$ $t$ 00
B.3 Antithetic Sampling
As shown by Albergo et al. [33], the loss function can become unstable around $t=0$ and $t=1$ for certain choices of $\gamma(t)$. To account for this, we implement antithetic sampling. This requires simultaneously computing the loss at both $x^{+}$ and $x^{-}$ where
@@ B.3 Antithetic Sampling
Both losses are computed using the same value of $z$ and subsequently averaged.
B.4 Diffusion Coefficient
An important inference-time parameter for models integrated with an SDE is the choice of $\epsilon(t)\geq 0$ which plays the role of a diffusion coefficient. Albergo et al. [33] note that the presence of $\gamma^{-1}(t)$ in the drift term seen in Fig. 5 can pose a numerical instability at the endpoints $t=0$ and $t=1$ during integration. For the choice $\epsilon_{\mathrm{const}}(t)=c$, they consider integrating from some nonzero time $t\gtrsim 0$ to $t\lesssim 1$ in order to avoid the singularity. Alternatively, one can design a form for $\epsilon(t)$ such that it vanishes at these endpoints. In OMatG, we opt for the latter approach and consider a diffusion coefficient, $\epsilon_{\mathrm{vanish}}(t)$, which vanishes at the endpoints
$\epsilon_{\mathrm{vanish}}(t)=\frac{c}{\left(1+e^{-\frac{t-\mu}{\sigma}}\right)\left(1+e^{-\frac{1-\mu-t}{\sigma}}\right)}.$ (9)
Here, $c$ dictates the magnitude of the diffusion, $\mu$ sets the times at which the midpoints between $\epsilon(t)=0$ and $\epsilon(t)=c$ are reached, and $\sigma$ controls the rate of this increase from $\epsilon(t)=0$ to $\epsilon(t)=c$. The only constraints on these parameters are that $c\geq 0$, $\mu>0$, and $\sigma>0$. Importantly, these parameters should be chosen such that they are near zero at the endpoints.
B.5 Interpolation with Periodic Boundary Conditions
We adopt a task-specific formulation for handling periodic boundary conditions with SIs tailored to flat tori, which are the relevant manifolds for fractional coordinates in crystal generation. We do not attempt to generalize stochastic interpolants (SIs) to arbitrary manifolds as in Riemannian flow matching [42].
As in FlowMM [32], in order to uniquely define the interpolating paths, we rely on shortest geodesic interpolation paths between pairs of fractional coordinates from $x_{0}$ and $x_{1}$, ensuring that interpolants are well-defined and differentiable. As noted in Section III.2.1, this shortest geodesic path is computed by first unwrapping one of the coordinates (say $x_{1}$) into its periodic image $x_{1}^{\prime}$, such that it is the closest image to $x_{0}$. We then compute, for example, the linear interpolant $x(t,x_{0},x_{1}^{\prime})=(1-t)x_{0}+tx_{1}^{\prime}$, as if in Euclidean space and finally wrap the interpolated path back onto the torus. (The geodesic is the same as the linear interpolant wrapped back into the box.)
We perform this procedure for all choices of interpolants. The reason for unwrapping according to the closest image first—for all interpolants—is because there are multiple ways to connect two points on a torus (e.g., in a periodic box one can connect two points with or without crossing the box boundaries). All periodic stochastic interpolants are then defined this way, by computing $x(t,x_{0},x_{1}^{\prime},z)=\alpha(t)x_{0}+\beta(t)x_{1}^{\prime}+\gamma(t)z$ in the unwrapped (Euclidean) space and wrapping back onto the torus. We emphasize that this procedure is important not only for the choice of interpolant, but also for the addition of the latent variable $\gamma(t)z$ which also moves the interpolation trajectory away from the geodesic. Our process yields exactly the same shortest-path geodesic as in FlowMM if using the linear interpolant, and thus recovers its corresponding conditional flow-matching loss. We depict our implementation of periodic stochastic interpolants in Fig. 6. We also demonstrate in Fig. 6c that averaging over the latent variable $\gamma(t)z$ recovers the deterministic base interpolant path, as required by the SI framework.
@@ B.5 Interpolation with Periodic Boundary Conditions
B.6 DFM Details
DFM allows for generative modeling of discrete sequences of tokens while respecting the discrete nature of the design space. As discussed, a parameterized neural network $p_{1|t}^{\theta}(x_{1}|x_{t})$ is learned, which attempts to predict the final sequence from the sequence at time $t$. Borrowing from Campbell et al. [44], we choose a conditional rate matrix $R_{t}(x_{t},i|x_{1})$—giving the rate of $x_{t}$ jumping to a different state $i$ given $x_{1}$—which generates the conditional flow $p_{t|1}(x_{t}|x_{1})$ of the form:
@@ B.6 DFM Details
In practice, Campbell et al. [44] show that we need not compute a full expectation, but rather, simply draw $x_{1}\sim p_{1|t}^{\theta}(x_{1}|x_{t})$, evaluate the conditional rate matrix $R_{t}^{\eta}(x_{t},i|x_{1})$, and perform an update of $x_{t}$ to $x_{t+\Delta t}$ with discrete time step $\Delta t$ directly from this by sampling $x_{t+\Delta t}$ according to
@@
B.7 Velocity Annealing
Velocity annealing—rescaling the learned velocity field during generation to increase velocity over time as $b^{\theta}(t,x)\rightarrow(1+st)\,b^{\theta}(t,x)$ with $s$ as an hyperparameter during integration—has been empirically shown to improve performance in a number of studies that apply CFM to physical systems [62, 63, 32]. For instance, Miller et al. [32] demonstrated that applying velocity annealing significantly improves performance in CSP and DNG benchmarks for materials. Motivated by these findings, we include velocity annealing in OMatG as a tunable hyperparameter, while emphasizing that this technique lacks a formal theoretical justification within the mathematical frameworks underlying flow models and stochastic interpolants.
B.8 Data-Dependent Coupling
SIs have been used with data-dependent couplings [64], where a coupling function $\nu(x_{0},x_{1})$ enables biasing of $x_{0}$ based on the sampled $x_{1}$. In OMatG, we incorporate an optional data-dependent coupling that enforces an ordering (i.e., a permutation on the order of atomic elements within a structure) that produces the minimum fractional-coordinate distance between each particle pair $(\bm{x}_{0}^{i},\bm{x}_{1}^{i})$ from structures $x_{0}\in\rho_{0}$ and $x_{1}\in\rho_{1}$. We find that the inclusion of this data-dependent coupling is optimal during hyperparameter tuning depending on the type of model: CSP models typically performed better without this coupling, but DNG models (see Tab. 11) can benefit in certain cases from minimizing traveled distance via permutation of elements.
Formally, our coupling is conditional on the sampled $(x_{0},x_{1})$ and is defined as